Find the steady-state temperature in a bar whose ends are located at x = 0 and
x = 15, if these ends are kept at 200°C and 100°C respectively. (Use Fourier series)
1
Expert's answer
2020-09-08T09:12:04-0400

There is the boundary value problem

"u_t(x, t) = ku_{xx}(x, t)" , for "t > 0" and "0 < x < L" (1)

"u(0, t) = T_0" and "u(L, t) = T_L" , for "t > 0"

"u(x, 0) = f(x)" , for "0 \leqslant x \leqslant L"

A steady-state temperature is one that does not depend on time.

Then ut = 0, so the heat equation (1) simplifies uxx = 0. Hence we are looking for a function us(x) defined for 0 < x < L such that

"\frac{∂^2u_s}{∂x^2}(x) = 0" , for "0 < x < L"

"u_s(0, t) = T_0" and "u_s(L,t) = T_L" for "t > 0"

The solution to this boundary value problem is easily found,since the general solution

of the differential equation is us(x) = Ax + B; where A and B are arbitrary constants.

Then the boundary conditions reduce to

us(0) = B = T0 and us(L) = AL + B = TL

We conclude that B = T0 and A = (TL – T0)/L, so the steady-state temperature is

"u_s(x) = (T_L – T_0)\frac{x}{L} + T_0"

"u_s(x) = (100 – 200)\frac{x}{15} + 200 = 200 – 6.66x"

Answer: 200 – 6.66x